# Neural Operator for Accelerating Coronal Magnetic Field Model

### v2

@@ 3. Experiments
All data used in the experiments are available from Hinode Science Data Centre Europe (https://sdc.uio.no/search/simulations). In our study, we will focus on the 3D magnetic data cube B along the $x$, $y$, and $z$ axes—denoted as $B_{x}$, $B_{y}$, and $B_{z}$, respectively. These components depict the 3D magnetic environment of the region of interest, in our case, an enhanced network area on the Sun. The data are located in the “en024048_hion”, a collection of 3D simulated magnetic data cubes generated using the Bifrost code to simulate a computational volume with a magnetic field topology similar to an enhanced network area on the Sun (Carlsson et al., 2016). We use the dataset of $B_{z}$ as one example to illustrate the data formats (the same applies to $B_{x}$ and $B_{y}$). The first snapshot of $B_{z}$ was captured at $t=3850$ seconds and the last snapshot of $B_{z}$ was captured at $t=5410$ seconds. Snapshots were recorded at 10-second intervals, resulting in 157 cubes. Therefore, $B_{z}$ dataset contains $\{B_{z}^{j}\}_{j=1}^{157}$. The size of each cube $B_{z}^{j}$ is $504\times 504\times 496$. The third dimension (i.e., 496) represents the height, which extends from the upper convection zone to the corona and is aimed at aiding the study of the solar chromosphere, a region that is challenging to model due to its complex dynamics and physics (Carlsson et al., 2016). The simulation volume measures 24 Mm $\times$ 24 Mm horizontally ($504\times 504$ pixels), with a depth of 2.4 Mm below the solar surface, and extends 14.4 Mm above, covering the upper part of the convection zone, the photosphere, chromosphere, transition region, and corona.
3.1. Data Prepossessing
@@
Given the constraints (limited memory space) of our GPU server, we downsampled the cube from $504\times 504\times 496$ to $504\times 504\times 100$ by selecting every five slices with index $1,6,11,\dots,496$. Our goal here is to learn the mapping from the bottom-most slice (the boundary condition of PDEs) to the remaining 99 slices as shown in Figure 1 (a). The deep learning method inputs the bottom-most image of size $504\times 504$ and outputs a 3D cube of size $504\times 504\times 99$. In addition, data normalization is implemented. 90% of the data is randomly selected for training.
3.2. Experiment Settings
In our experiments, we implemented three other deep learning methods to compare with TFNO.
3.2.1. TNFO Structure
The code (https://github.com/neuraloperator/neuraloperator) we used is publicly available. To increase the complexity of the neural network, and thus its expressive power, we set the number of modes retained in each dimension in the Fourier layer to $(64,64)$ and set the hidden channel to $64$, thus enabling the model to efficiently address the complex spatial structure inherent in the data. The model uses the Tucker decomposition with a rank$=0.42$, which balances expressiveness with computational and storage efficiency. It captures the most important interactions between input data elements while discarding less important ones to create a more compact model. In addition to the basic Lifting Layer, FNO Blocks, and Projection Layer mentioned in Section 2.2, a layer of Multi-Layer Perceptrons (MLPs) is added after the FNO Blocks to output the desired size of the 3D cube ($504\times 504\times 99$) since the original code can only output 2D image. Additionally, it enables an additional non-linear processing stage, improving its capability to model complex functions and enhancing its ability to learn a diverse array of PDE solutions.
3.2.2. Vision Transformer (ViT)
The ViT (Dosovitskiy et al., 2020) used in our paper is a modification of the transformer architecture traditionally used for Natural Language Processing (NLP). Our ViT is adapted to process the input image in the form of patches, which are embedded with masks and then passed through the transformer encoder. The model processes $504\times 504$ images using $8\times 8$ patches. With 128-dimensional embedding and a multi-head attention mechanism spanning 2 layers of 8 heads. A learnable mask embedding is added to the patch embedding sequence, and positional embedding is added to the spatial information lost during the patch embedding process. After converter processing, the output is summarized and decoded to the desired number of output channels for our prediction.
3.2.3. CNN-RNN and CNN-LSTM
We use two different encoder-decoder structures: The CNN-RNN (Wang et al., 2016) and CNN-LSTM (Hara et al., 2018). We use Convolutional Neural Networks (CNN) for feature extraction and Recurrent Neural Networks (RNN) or Long Short-Term Memory Networks (LSTM) for modeling sequential images. The CNN component consists of three convolutional layers that progressively increase the number of feature channels from 1 to 256 while down-sampling the image dimensions by a factor of 2 after each layer. This is done via convolution with a stride of 2, effectively reducing the spatial dimensions and extracting higher-level features. In CNN-RNN, the main component of the recurrent network is a series of gated recurrent units (GRUs), which are chosen for their ability to efficiently model time series. The hidden state size of the GRU units in this model is 256, which reflects a balance between model complexity and computational efficiency, allowing the network to capture a large amount of temporal information without being too computationally demanding. Meanwhile, the CNN-LSTM uses LSTM cells with a hidden state size of 256, which is consistent with the output dimension of the CNN.
3.2.4. Optimizer and Scheduler
All deep learning methods are trained using the Adam optimizer (Kingma and Ba, 2014) and the ReduceLROnPlateau scheduler with 300 epochs and a batch size of 10. An Adam optimizer was used with an initial learning rate of 0.001 and a weight decay of 0.0001. The ReduceLROnPlateau scheduler was used to adjust the learning rate according to the performance of the model, with a learning rate reduction factor of 0.8 and a patience of 2. Huber Loss (Taylor et al., 2023) is used for training for all the methods to approximate PDE solutions. It is a hybrid loss function, known for blending the properties of mean squared error and mean absolute error, demonstrating robustness against outliers—a feature when dealing with the intrinsic discontinuities and sharp gradients characteristic of PDEs.
3.3. Results
3.3.1. Evaluation from an AI/ML Perspective
@@
Figure 2. Training and test error curves for different epochs of TFNO, ViT, CNN-RNN, and CNN-LSTM.
[…]
Example.
Figure 2 shows the training error, i.e., Huber Loss, and test error, i.e., Mean Squared Error (MSE), over 300 epochs for the TFNO, ViT, CNN-RNN, and CNN-LSTM. The training error of TFNO and ViT shows an overall decreasing trend and converges to a much lower error than CNN-RNN and CNN-LSTM. Meanwhile, TFNO shows a very stable and smooth training error compared to all other deep learning methods. In addition, we observe a divergence in the generalization capabilities between the ViT and the TFNO. TFNO shows a lower training error before 220 epochs compared to ViT while ViT shows a slightly lower training error after 220 epochs. However, this improvement does not translate equivalently to the test phase, where ViT has a significantly higher test error compared to TFNO. This is caused by overfitting of the ViT, a condition where a model’s excessive complexity captures not only the underlying data distribution but also the noise within the training set. This is potentially caused by ViT’s parameter-rich architecture that encompasses the largest number of model parameters among all deep learning methods. Conversely, TFNO has consistently better test performance, suggesting a more robust generalization capacity. This pattern highlights the potential advantages of TFNO to mitigate overfitting. While CNN-RNN and CNN-LSTM have similar training errors, CNN-RNN outperforms CNN-LSTM in test error. However, both methods have much higher test errors compared to TFNO and ViT.
@@
Table 1. Test performance comparison of TFNO, ViT, CNN-RNN, and CNN-LSTM for B in terms of Average Training Time (one epoch), MSE, $R^{2}$, RE, and MAE.
Model Time $\downarrow$ MSE $\downarrow$ $R^{2}\uparrow$ RE $\downarrow$ MAE $\downarrow$
TFNO 31.96s 0.0502 0.9498 0.2205 0.0831
ViT 39.25s 0.1240 0.8759 0.3498 0.1417
CNN-RNN 32.93s 0.2514 0.7485 0.4995 0.2230
CNN-LSTM 37.10s 0.3257 0.6743 0.5686 0.2940
@@
Figure 3. The visualization of prediction from TFNO, ViT, CNN-RNN, and CNN-LSTM, compared with ground truth at different heights: (a) $B_{x}$; (b) $B_{y}$; (c) $B_{z}$.
[…]
Example.
@@
Figure 4. The visualization of error maps from TFNO, ViT, CNN-RNN, and CNN-LSTM at different heights: (a) $B_{x}$; (b) $B_{y}$; (c) $B_{z}$.
[…]
Example.
@@
In addition to MSE, Table 1 summarizes the performance of different deep learning methods in terms of Average training time of one epoch, R-square ($R^{2}$), Relative Error (RE), and Mean Absolute Error (MAE). TFNO shows the best test performance among all deep learning methods. TFNO has the lowest MSE, RE, and MAE. TFNO achieves $R^{2}=0.9498$ while the ViT achieves the second-best $R^{2}=0.8759$. This verifies that ViT has a suboptimal generalization capability indicated by the lowest training error but a higher test error compared to TFNO. On the other hand, the CNN-RNN and CNN-LSTM (i.e., CNN-based recurrent architectures) show the second-worst and worst test performance, respectively, which are not comparable with TFNO and ViT. Meanwhile, they do not offer computational advantages over ViT as reflected in their average training time for each epoch. Overall, TFNO not only shows the best test performance but also shows the best training time efficiency.
@@
Figure 3 shows the visualization of prediction from TFNO, ViT, CNN-RNN, and CNN-LSTM, compared with ground truth at different heights. In Figures 3a, 3b, and 3c, the selected slices are captured at different heights of the cube. The first column in each Figure is the input data of the deep learning methods. The second through fifth columns correspond to heights of approximately 0 Mm, 2 Mm, 4 Mm, 6 Mm, and 10 Mm, respectively. These slices were chosen to demonstrate the diversity of magnetic field strengths at different heights. The first row represents the “ground truth”, which serves as a benchmark for evaluating the subsequent rows – TFNO, ViT, CNN-RNN, and CNN-LSTM.
From these three figures, we can observe that the prediction results of TFNO and ViT are almost the same as the ground truth while CNN-RNN and CNN-LSTM reveal remarkable discrepancies in fine details when compared with the ground truth at heights =0, 2 Mm, 4 Mm, 6 Mm, and 10 Mm. Specifically, CNN-RNN and CNN-LSTM cannot learn the pattern of the ground truth at heights = 0 Mm.
@@
To better visualize the predictions of different deep learning methods, Figure 4 shows the error maps of $B_{x}$, $B_{y}$, and $B_{z}$ from TFNO, ViT, CNN-RNN, and CNN-LSTM at different heights of approximately 0 Mm, 2 Mm, 4 Mm, 6 Mm, and 10 Mm. These different heights are corresponding to the first, second, third, fourth, and fifth columns, respectively. These three figures are obtained by subtracting the predicted results from the ground truth and then taking the absolute value. The blue color represents a small error, indicating that the predicted values are closer to the actual values. These regions can be interpreted as areas where the deep learning method has reached a high level of accuracy, successfully capturing the underlying patterns and dynamics of the data. Conversely, the red color represents a large error, indicating an error between the output from the deep learning method and the ground truth. The darkness of the red color is proportional to the magnitude of the error, with darker red indicating a larger error.
@@
Figure 4 shows that there are prediction errors for all deep learning methods around a height of approximately 0 Mm. However, TFNO and ViT still perform much better compared to CNN-RNN and CNN-LSTM. This observation can be attributed to several potential factors, including the inherent complexity of the data. While it may be challenging to find differences between TFNO and ViT in Figures 3a, 3b, and 3c, Figures 4a, 4b, and 4c allow us to observe that TFNO only has the smallest prediction errors at heights of approximately 2 Mm, 4 Mm, 6 Mm, and 10 Mm, where the error map is mostly dark blue (i.e., no errors). ViT shows serpentine, thin-line errors, indicating a deficiency in capturing fine details as compared to TFNO. The CNN-RNN and CNN-LSTM exhibit learning capabilities but present significantly larger errors than TFNO and ViT, which are not only quantitatively substantial but also lack meaningful physical interpretation.
@@
In conclusion, TFNO stands out for its robustness and ability to capture detailed features with the lowest test error, indicating its potential suitability for complex tasks requiring high precision. In contrast, while ViT shows promise in certain aspects, its performance is shadowed by its shortcomings in complex structure predictions.
@@
Figure 5. 2D histogram of TFNO, ViT, CNN-RNN, and CNN-LSTM to the ground truth, at height = 2.0 Mm, 7.0 Mm, and 11.0 Mm. Color gradient denotes the number of data points.
[…]
Example.
3.3.2. Evaluation from a Physics Perspective
@@
Figure 5 shows the 2D histograms that compare predictions from different deep learning methods (TFNO, ViT, CNN-RNN, and CNN-LSTM) to the ground truth across various height = 2.0 Mm, 7.0 Mm, and 11.0 Mm. Each method demonstrates a scattered distribution of points along the diagonal, indicating a linear correlation between predicted and true magnetic field strengths. At a height of 2.0 Mm, closer to the surface, the TFNO and ViT show a denser clustering of points around the diagonal, a pattern less concentrated in the CNN-LSTM and CNN-RNN. At greater heights of 7.0 Mm and 11.0 Mm, TFNO and ViT consistently keep points near the diagonal, showcasing sustained prediction accuracy. On the contrary, CNN-LSTM and CNN-RNN exhibit a wider spread of points, hinting at a decline in predictive precision.
@@
Figure 6. Comparison of physics quantities derived from TFNO, ViT, CNN-RNN, and CNN-LSTM. Upper panel: Variation of magnetic field strength as a function of height. Lower panel: Variation of current density as a function of height.
[…]
Example.
@@
Moreover, we have evaluated physics-based measurements varying with height to assess each model’s ability to reproduce the MHD model outputs. Figure 6 displays magnetic field strength (B0) and current density ($|J|$) reveals that, above a height of 2 Mm, all models produce patterns that resemble the ground truth, indicative of a stable environment with low magnetic field strength. Below this height, only the TFNO model’s predictions align closely with the actual measurements. Similarly, with current density, TFNO maintains a near-ground truth trend below a height of 1.5 Mm. At a height above 1.5 Mm, characterized by low current, all deep learning methods follow a similar trajectory. Overall, TFNO delivers the most consistent and accurate predictions of these physical quantities, particularly close to the surface where the magnetic settings are complex.
@@
Figure 7. Comparison of field orientation derived from TFNO, ViT, CNN-RNN, and CNN-LSTM. Upper panel: Variation of inclination angle as a function of height. Lower panel: Variation of azimuthal angle as a function of height.
[…]
Example.
@@
Figure 7 illustrates the variation of the horizontally averaged inclination angle $\theta$ and azimuthal angle $\phi$ as a function of height, derived from different deep learning methods compared to the ground truth. Please note that a binary map has been applied to focus solely on magnetic field strengths that exceed 2.5 times the background level. This approach helps to disregard weaker fields, which tend to be noisier and less significant, while still taking into account their potential influence on the calculations of field orientation metrics. All deep learning methods can reproduce the inclination angles well after height $>$ 1Mm. At a lower height below 1 Mm, ViT outperforms other models regarding the inclination angles. The azimuthal angle variation, however, demonstrates a significant divergence between the deep learning methods and the ground truth at the lower height (below height = 2 Mm). TFNO and ViT appear to align with the ground truth, but TFNO converges faster than ViT. Although CNN-LSTM and CNN-RNN follow the general trend, they do not closely align with the ground truth. At the higher layer (height $>$ 12 Mm), TFNO continues to track the ground truth, whereas ViT starts to diverge. The inclination and azimuthal angles are important for understanding the orientation of the magnetic field. The consistency of the TFNO with the ground truth suggests it may provide a more physically accurate magnetic field orientation with height, compared with other deep learning methods.
4. Conclusion

### v3

@@ IV Experiments and Results
All data used in the experiments are available from Hinode Science Data Centre Europe (https://sdc.uio.no/search/simulations). In our study, we will focus on the 3D magnetic data cube B along the $x$, $y$, and $z$ axes—denoted as $B_{x}$, $B_{y}$, and $B_{z}$, respectively. These components depict the 3D magnetic environment of the region of interest, in our case, an enhanced network area on the Sun. The data are located in the “en024048_hion”, a collection of 3D simulated magnetic data cubes generated using the Bifrost code to simulate a computational volume with a magnetic field topology similar to an enhanced network area on the Sun [51]. We use the dataset of $B_{z}$ as one example to illustrate the data formats (the same applies to $B_{x}$ and $B_{y}$). The first snapshot of $B_{z}$ was captured at $t=3850$ seconds and the last snapshot of $B_{z}$ was captured at $t=5410$ seconds. Snapshots were recorded at 10-second intervals, resulting in 157 cubes. Therefore, $B_{z}$ dataset contains $\{B_{z}^{j}\}_{j=1}^{157}$. The size of each cube $B_{z}^{j}$ is $504\times 504\times 496$. The third dimension (i.e., 496) represents the height, which extends from the upper convection zone to the corona and is aimed at aiding the study of the solar chromosphere, a region that is challenging to model due to its complex dynamics and physics [16]. The simulation volume measures 24 Mm $\times$ 24 Mm horizontally ($504\times 504$ pixels), with a depth of 2.4 Mm below the solar surface, and extends 14.4 Mm above, covering the upper part of the convection zone, the photosphere, chromosphere, transition region, and corona.
IV-A Data Preprocessing
@@
Given the constraints (limited memory space) of our GPU server, we downsampled the cube from $504\times 504\times 496$ to $504\times 504\times 100$ by selecting every five slices with index $1,6,11,\dots,496$. Our goal here is to learn the mapping from the bottom-most slice (the boundary condition of PDEs) to the remaining 99 slices as shown in Fig. 1 (a). The deep learning method inputs the bottom-most image of size $504\times 504$ and outputs a 3D cube of size $504\times 504\times 99$. In addition, data normalization is implemented. 90% of the data is randomly selected for training.
@@
Figure 2: Training and test error curves for different epochs of GL-FNO, ViT, CNN-RNN, and CNN-LSTM.
IV-B Experiment Settings
In our experiments, we implemented three other deep learning methods together with GL-FNO.
IV-B1 GL-FNO Structure
Our GL-FNO is built based on publicly available code (https://github.com/neuraloperator/neuraloperator). For the local branch, we set the number of modes retained in each dimension in the Fourier layer to $(64,64)$. We set the hidden channel to $128$, thus enabling the model to efficiently address the complex spatial structure inherent in the data. The model uses the Tucker decomposition with a rank$=0.5$, which balances expressiveness with computational and storage efficiency. It captures the most important interactions between input data elements while discarding less important ones to create a more compact model. In addition to the basic Lifting Layer, FNO Blocks, and Projection Layer mentioned in Section III, a layer of Multi-Layer Perceptrons (MLPs) is added after the FNO Blocks to output the desired size of the 3D cube ($504\times 504\times 99$) since the original code can only output 2D image. Additionally, it enables an additional non-linear processing stage, improving its capability to model complex functions and enhancing its ability to learn a diverse array of PDE solutions. We keep the same structure for the global branch as the local one.
IV-B2 Vision Transformer (ViT)
The ViT [42] used in our paper is a modification of the transformer architecture traditionally used for Natural Language Processing (NLP). Our ViT is adapted to process the input image in the form of patches, which are embedded with masks and then passed through the transformer encoder. The model processes $504\times 504$ images using $8\times 8$ patches. With 128-dimensional embedding and a multi-head attention mechanism spanning 2 layers of 8 heads. A learnable mask embedding is added to the patch embedding sequence, and positional embedding is added to the spatial information lost during the patch embedding process. After converter processing, the output is summarized and decoded to the desired number of output channels for our prediction.
IV-B3 CNN-RNN and CNN-LSTM
We use two different encoder-decoder structures: The CNN-RNN [43] and CNN-LSTM [44]. We use Convolutional Neural Networks (CNN) for feature extraction and Recurrent Neural Networks (RNN) or Long Short-Term Memory Networks (LSTM) for modeling sequential images. The CNN component consists of three convolutional layers that progressively increase the number of feature channels from 1 to 256 while down-sampling the image dimensions by a factor of 2 after each layer. This is done via convolution with a stride of 2, effectively reducing the spatial dimensions and extracting higher-level features. In CNN-RNN, the main component of the recurrent network is a series of gated recurrent units (GRUs), which are chosen for their ability to efficiently model time series. The hidden state size of the GRU units in this model is 256, which reflects a balance between model complexity and computational efficiency, allowing the network to capture a large amount of temporal information without being too computationally demanding. Meanwhile, the CNN-LSTM uses LSTM cells with a hidden state size of 256, which is consistent with the output dimension of the CNN.
IV-B4 Optimizer and Scheduler
All deep learning methods are trained using the Adam optimizer [52] and the ReduceLROnPlateau scheduler with 300 epochs and a batch size of 10. An Adam optimizer was used with an initial learning rate of 0.001 and a weight decay of 0.0001. The ReduceLROnPlateau scheduler was used to adjust the learning rate according to the performance of the model, with a learning rate reduction factor of 0.8 and patience of 2. Huber Loss [53] is used for training for all the methods to approximate PDE solutions. It is a hybrid loss function, known for blending the properties of mean squared error and mean absolute error, demonstrating robustness against outliers—a feature when dealing with the intrinsic discontinuities and sharp gradients characteristic of PDEs.
IV-C Results
We mainly aim to address two questions
  • Section IV-C1: What is the performance of GL-FNO compared with other deep learning methods?
  • Section IV-C2: Do the solutions (i.e., prediction) generated from GL-FNO make sense from the perspective of a solar physicist?
IV-C1 Evaluation from an AI/ML Perspective
Fig. 2 shows the training error, i.e., Huber Loss, and test error, i.e., Mean Squared Error (MSE), over 300 epochs for the GL-FNO, ViT, CNN-RNN, and CNN-LSTM. The training and test errors of GL-FNO and ViT show an overall decreasing trend and converge to a much lower error than CNN-RNN and CNN-LSTM. GL-FNO not only converges to the lowest training and test error but also shows very stable and smooth training and test error curves compared to all other deep learning methods.
@@ IV-C1 Evaluation from an AI/ML Perspective
TABLE I: Test performance comparison of GL-FNO, ViT, CNN-RNN, and CNN-LSTM for B in terms of MSE, $R^{2}$, RE, MAE, PSNR, SSIM.
Model MSE $\downarrow$ $R^{2}\uparrow$ RE $\downarrow$ MAE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
GL-FNO 0.0331 0.9669 0.1756 0.0672 50.58 0.9883
ViT 0.1240 0.8759 0.3498 0.1417 35.42 0.8856
CNN-RNN 0.2514 0.7485 0.4995 0.2230 30.72 0.7781
CNN-LSTM 0.3257 0.6743 0.5686 0.2940 28.91 0.7356
@@
Figure 3: The visualization of prediction from GL-FNO, ViT, CNN-RNN, and CNN-LSTM, compared with ground truth at different heights: (a) $B_{x}$; (b) $B_{y}$; (c) $B_{z}$.
@@
Figure 4: The visualization of error maps from GL-FNO, ViT, CNN-RNN, and CNN-LSTM at different heights: (a) $B_{x}$; (b) $B_{y}$; (c) $B_{z}$.
@@
In addition to MSE, Table I summarizes the performance of different deep learning methods in terms of R-square ($R^{2}$), Relative Error (RE), Mean Absolute Error (MAE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index Measure (SSIM). GL-FNO shows the best test performance among all deep learning methods. GL-FNO has the lowest MSE, RE, and MAE and the highest $R^{2}$, PSNR, and SSIM. GL-FNO achieves $R^{2}=0.9498$, PSNR = 50.58, and SSIM = 0.9883 while the ViT achieves the second-best $R^{2}=0.8759$, PSNR = 35.42, and SSIM = 0.8856. On the other hand, the CNN-RNN and CNN-LSTM (i.e., CNN-based recurrent architectures) show the second-worst and worst test performance, respectively, which are not comparable with GL-FNO and ViT. Overall, GL-FNO shows the best test performance. In our Appendix Experiments with Different Neural Operators, we also compared different neural operators with our GL-FNO, where the same conclusion can be obtained.
@@
Fig. 3 shows the visualization of prediction from GL-FNO, ViT, CNN-RNN, and CNN-LSTM, compared with ground truth at different heights. In Figs 3a, 3b, and 3c, the selected slices are captured at different heights of the cube. The first column in each Figure is the input data of the deep learning methods. The second through fifth columns correspond to heights of approximately 0 Mm, 2 Mm, 4 Mm, 6 Mm, and 10 Mm, respectively. These slices were chosen to demonstrate the diversity of magnetic field strengths at different heights. The first row represents the “ground truth”, which serves as a benchmark for evaluating the subsequent rows – GL-FNO, ViT, CNN-RNN, and CNN-LSTM. From these three figures, we can observe that the prediction results of GL-FNO and ViT are almost the same as the ground truth while CNN-RNN and CNN-LSTM reveal remarkable discrepancies in fine details when compared with the ground truth at heights = 0, 2 Mm, 4 Mm, 6 Mm, and 10 Mm. Specifically, CNN-RNN and CNN-LSTM cannot learn the pattern of the ground truth at heights = 0 Mm.
@@
To better visualize the predictions of different deep learning methods, Fig. 4 shows the error maps of $B_{x}$, $B_{y}$, and $B_{z}$ from GL-FNO, ViT, CNN-RNN, and CNN-LSTM at different heights of approximately 0 Mm, 2 Mm, 4 Mm, 6 Mm, and 10 Mm. These different heights are corresponding to the first, second, third, fourth, and fifth columns, respectively. These three figures are obtained by subtracting the predicted results from the ground truth and then taking the absolute value. The blue color represents a small error, indicating that the predicted values are closer to the actual values. These regions can be interpreted as areas where the deep learning method has reached a high level of accuracy, successfully capturing the underlying patterns and dynamics of the data. Conversely, the red color represents a large error, indicating an error between the output from the deep learning method and the ground truth. The darkness of the red color is proportional to the magnitude of the error, with darker red indicating a larger error.
@@
Fig. 4 shows that there are prediction errors for all deep learning methods around a height of approximately 0 Mm. However, GL-FNO and ViT still perform much better compared to CNN-RNN and CNN-LSTM. This observation can be attributed to several potential factors, including the inherent complexity of the data. While it may be challenging to find differences between GL-FNO and ViT in Figs 3a, 3b, and 3c, Figs 4a, 4b, and 4c allow us to observe that GL-FNO only has the smallest prediction errors at heights of approximately 2 Mm, 4 Mm, 6 Mm, and 10 Mm, where the error map is mostly dark blue (i.e., no errors). ViT shows serpentine, thin-line errors, indicating a deficiency in capturing fine details as compared to GL-FNO. The CNN-RNN and CNN-LSTM exhibit learning capabilities but present significantly larger errors than GL-FNO and ViT, which are not only quantitatively substantial but also lack meaningful physical interpretation.
@@
In conclusion, GL-FNO stands out for its robustness and ability to capture detailed features with the lowest test error, indicating its potential suitability for complex tasks requiring high precision. In contrast, while ViT shows promise in certain aspects, its performance is shadowed by its shortcomings in complex structure predictions.
@@
Figure 5: 2D histogram of GL-FNO, ViT, CNN-RNN, and CNN-LSTM to the ground truth, at height = 2.0 Mm, 7.0 Mm, and 11.0 Mm. Color gradient denotes the number of data points.
IV-C2 Evaluation from a Physics Perspective
@@
Fig. 5 shows the 2D histograms that compare predictions from different deep learning methods (GL-FNO, ViT, CNN-RNN, and CNN-LSTM) to the ground truth across various height = 2.0 Mm, 7.0 Mm, and 11.0 Mm. Each method demonstrates a scattered distribution of points along the diagonal, indicating a linear correlation between predicted and true magnetic field strengths. At a height of 2.0 Mm, closer to the surface, the GL-FNO and ViT show a denser clustering of points around the diagonal, a pattern less concentrated in the CNN-LSTM and CNN-RNN. At greater heights of 7.0 Mm and 11.0 Mm, GL-FNO consistently keeps points near the diagonal, showcasing sustained prediction accuracy. On the contrary, ViT, CNN-LSTM and CNN-RNN exhibit a wider spread of points, hinting at a decline in predictive precision.
@@
Figure 6: Comparison of physics quantities derived from GL-FNO, ViT, CNN-RNN, and CNN-LSTM. Upper panel: Variation of magnetic field strength as a function of height. Lower panel: Variation of current density as a function of height.
@@
Moreover, we have evaluated physics-based measurements varying with height to assess each model’s ability to reproduce the MHD model outputs. Fig. 6 displays magnetic field strength ($|B|$) and current density ($|J|$) revealing that, above a height of 2 Mm, all models produce patterns that resemble the ground truth, indicative of a stable environment with low magnetic field strength. Below this height, only the GL-FNO model’s predictions align closely with the actual measurements. Similarly, with current density, GL-FNO maintains a near-ground truth trend below a height of 1.5 Mm. At a height above 1.5 Mm, characterized by low current, all deep learning methods follow a similar trajectory. Overall, GL-FNO delivers the most consistent and accurate predictions of these physical quantities, particularly close to the surface where the magnetic settings are complex.
@@
Figure 7: Comparison of field orientation derived from GL-FNO, ViT, CNN-RNN, and CNN-LSTM. Upper panel: Variation of inclination angle as a function of height. Lower panel: Variation of azimuthal angle as a function of height.
@@
Fig. 7 illustrates the variation of the horizontally averaged inclination angle $\theta$ and azimuthal angle $\phi$ as a function of height, derived from different deep learning methods compared to the ground truth. Please note that a binary map has been applied to focus solely on magnetic field strengths that exceed 2.5 times the background level. This approach helps to disregard weaker fields, which tend to be noisier and less significant, while still taking into account their potential influence on the calculations of field orientation metrics. All deep learning methods can reproduce the inclination angles well after height $>$ 1 Mm. At a lower height below 1 Mm, ViT outperforms other models regarding the inclination angles. The azimuthal angle variation, however, demonstrates a significant divergence between the deep learning methods and the ground truth at the lower height (below height = 2 Mm). GL-FNO and ViT appear to align with the ground truth, but GL-FNO converges faster than ViT. Although CNN-LSTM and CNN-RNN follow the general trend, they do not closely align with the ground truth. At the higher layer (height $>$ 13 Mm), GL-FNO continues to track the ground truth, whereas ViT starts to diverge. The inclination and azimuthal angles are important for understanding the orientation of the magnetic field. The consistency of the GL-FNO with the ground truth suggests it may provide a more physically accurate magnetic field orientation with height, compared with other deep learning methods.
V Conclusion